# Stabilizing Fractional Dynamical Networks Suppresses Epileptic Seizures

### v1

@@ Discussion
Our stabilizing control framework successfully stabilized 77% of initially unstable seizures and achieved approximately 49% amplitude reduction across all cases. The control produced a similar average amplitude reduction in both initially stable and unstable seizures, showcasing effective suppression of ictal activity regardless of mathematical stability status. Patient HUP78, who developed epilepsy from traumatic injury, failed to achieve stabilization in all five of their seizures, including three of their seizures that were initially stable. This complete failure suggests that traumatic brain injury-induced structural damage, gliosis, and altered network connectivity [39] may require control strategies different from other epilepsy etiologies. Furthermore, the eight seizures that failed to stabilize exhibited severely ill-conditioned optimization problems, with the median condition numbers 50x larger than those of successful cases (1383 vs 28), which suggests that a feasible solution to the optimization problem presented in (12) does not exist. Physiologically, this may indicate that certain seizure dynamics involve such profound network reorganization or high-dimensional complexity that simple linear coupling modifications cannot restore stability [40]. Importantly, most control failures occurred in seizures with good model fits, indicating that accurate system identification does not guarantee a feasible solution to the optimization problem in (12). Future work will focus on developing conditions to guarantee a feasible solution by verifying stabilizability with state-feedback.
Certain network configurations that are well-characterized by fractional-order models may be fundamentally resistant to state-feedback interventions due to structural constraints, pathological states, or intrinsic nonlinearities that emerge during control attempts [40]. Patients with structural lesions, certain epilepsy etiologies, or particular network architectures may require alternative control strategies, such as nonlinear control, higher-order perturbations, or targeting different network nodes. Larger studies grouping patients by clinical characteristics may reveal which characteristics are more favorable to state-feedback control versus those requiring more sophisticated interventions [41, 40].
Fractional-Order Framework Complements Existing Epilepsy Modeling and Control Approaches
Our fractional-order system approach extends existing methods for characterizing and controlling epileptic seizures. Network fragility approaches [5, 29] use linear time-varying models to quantify network sensitivity to perturbations, identifying nodes where small synaptic changes can destabilize the network. Neural fragility achieved 76% accuracy in predicting surgical outcomes by identifying fragile regions [42]. Our fractional-order exponents ($\alpha$) quantify history-dependence in neural dynamics, a property that integer-order models cannot capture. While fragility focuses on instantaneous network sensitivity, $\alpha$ reveals how strongly current brain states depend on past activity across multiple timescales.
Other modeling approaches include nonlinear dynamical analysis methods that describe seizures as transitions between chaotic and ordered states [43]. These models require extensive neuroimaging and computational resources, whereas fractional-order modeling achieves an interpretable characterization using iEEG data with efficient parameter estimation. One such example is the Virtual Epileptic Patient using the Epileptor model, which characterizes seizure bifurcations through fast-slow nonlinear dynamics [44].
Responsive neurostimulation (RNS) achieved 37.9% seizure reduction during initial 12-week trials, improving to 53% at two years and 75% at nine years [45, 46, 47]. Vagus nerve stimulation showed 40-50% response rates with 5-10% achieving seizure freedom [48]. Our stabilizing control achieved 77% stabilization success in initially unstable seizures with 48% average amplitude reduction. Direct comparison is limited by study design differences, with our results representing in-silico open-loop control versus implanted in vivo closed-loop devices.
@@
Several limitations warrant consideration. Our small sample size (10 patients, 35 seizures) limited statistical power despite large effect sizes, and all patients had focal epilepsy, restricting generalizability to other epilepsy types. Model quality varied across segments, with post-ictal periods being particularly challenging, suggesting certain brain states may violate linear fractional-order assumptions. Our control method could face real-world implementation challenges, including robustness, sensing delays, stimulation artifacts, and safety constraints. Additionally, we focused on eigenvalues and fractional-order exponents ($\alpha$), but did not analyze eigenvectors. They could identify the spatial patterns of network reorganization across the seizure states and reveal which electrode drives transitions between states, which has the potential to improve targeting for neurostimulation. We focused exclusively on network dynamics without incorporating clinical features such as lesion characteristics, medications, or comorbidities that may influence seizure evolution.
Larger cohorts would enable the identification of patient subgroups with distinct seizure dynamics profiles and improve statistical power for detecting group-level patterns. A systematic investigation of window size and stride parameters could optimize the temporal resolution for capturing seizure transitions. Extending the estimation framework to better capture post-ictal dynamics that potentially violate linear assumptions in our analysis may improve our understanding of post-seizure recovery. Finally, our control framework was only validated through in silico simulations using estimated models from iEEG data, not in real brain tissue. Real-world implementation faces model accuracy, unmodeled dynamics, hardware delays, artifacts, and safety constraints that cannot be fully captured in computational models. Closed-loop implementation and validation in an animal model or clinical setting would assess the real-world feasibility and efficacy of fractional-order model-based neurostimulation for seizure suppression.
@@ Methods
where ${x}[k]=(x_{1}[k],x_{2}[k],\ldots,x_{N}[k])^{{T}}$ is a state vector of the $N$ pre-processed iEEG recordings at discrete time step $k\in\mathbb{N}$. $A\in\mathbb{R}^{N\times N}$ is a matrix that models the spatial coupling between states. The multi-scale behavior is described using a discrete-time fractional derivative operator $\Delta^{\alpha}$, which is the Grünwald-Letnikov discretization of the fractional derivative (Chpt. 2, [17]). The vector of fractional-order exponents ${\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{N})^{\mathsf{T}}$ computes the multi-scale property for each recording, which determines how strongly each measurement depends on its past measurements. Smaller values of $\alpha$ indicate a stronger long-range temporal dependency. A value of $\alpha=1$ corresponds to a standard first-order linear time-invariant dynamical system with no multi-scale behavior. This relationship is described for each electrode $i$ by:
@@
To improve the estimation of $A$ and account for potential unmeasured influences, we used an iterative estimation approach [52] to estimate the parameters and unknown inputs. We augmented Eq. (1) to include an input term:
@@
where $u[k]$ represents any external inputs to the system from unmeasured sources (e.g., brain regions outside the recording area, subcortical inputs, or modeling errors). The matrix $B$ was initialized heuristically from $A^{(0)}$, which is the estimated $A$ matrix during the first iteration, by selecting columns that represent the most influential electrodes in the network, assuming half the recorded electrodes could receive external inputs. For each time iteration $i$, we estimated the inputs by solving a least absolute shrinkage and selection operator (LASSO) problem:
@@
The model quality was evaluated by simulating signals using the estimated parameters $A$ and $\alpha$. Starting from Eq. (1) and expanding the fractional derivative using the Grünwald-Letnikov weights from Eq. (3) with $\Psi(\alpha_{i},0)=1$:
@@
This controller was applied to ictal windows, starting at seizure onset. We evaluated controller effectiveness by comparing the eigenvalues of $A_{0}$ matrix (uncontrolled) and $A_{0}+\tilde{A}$ matrix (controlled) during the ictal period. Additionally, to visualize the effect of control on signal dynamics, we reconstructed signals using Eq. (9), substituting $A_{\text{ctrl}}=A+\tilde{A}$ for $A$, to visualize the effect of control on signal amplitudes.

### v2

@@ Discussion
Patient HUP78, who developed epilepsy from traumatic injury, failed to achieve stabilization in all five of their seizures, including three of their seizures that were initially stable. This complete failure suggests that traumatic brain injury-induced structural damage, gliosis, and altered network connectivity [61] may require control strategies different from other epilepsy etiologies. Additionally, the eight seizures that failed to stabilize exhibited severely ill-conditioned optimization problems, with the median condition numbers 50x larger than those of successful cases (1383 vs 28), which suggests that a feasible solution to the optimization problem presented in (12) may not exist. Physiologically, this may indicate that certain seizure dynamics involve such profound network reorganization or high-dimensional complexity that linear coupling modifications may be unable to restore stability [62]. Importantly, most control failures occurred in seizures with good model fits, indicating that accurate system identification does not guarantee a feasible solution to the stabilizing control problem formulated as an optimization problem in (12). Future work will focus on developing feasibility conditions for solving the stabilizing control problem.
@@
A potential limitation of this study is the sample size (10 patients, 35 seizures), which may limit statistical power despite large effect sizes. Larger cohorts would enable the identification of patient subgroups with distinct seizure dynamics profiles and improve statistical power for detecting group-level patterns.
We observed that the goodness of fit of our model varied across segments, with post-ictal periods being particularly challenging. In our work, we considered a 3 second sliding window to estimate fractional-order dynamical networks. In future work, we will investigate the effects of window size selection on estimation. A systematic investigation of window size and stride parameters could optimize the temporal resolution for capturing seizure transitions. Additionally, more work is needed to assess the sensitivity of our estimation scheme to noise and extreme perturbations.
Our current control method may face real-world implementation challenges. In future work, we plan to address the issue of robustness and develop a scheme to overcome delays. Furthermore, our current strategy ignores stimulation artifacts and safety constraints, which we plan to address in a future study.
In this study, we focused on analyzing eigenvalues and fractional-order exponents ($\alpha$) of fractional-order dynamical networks, which show significant promise. In future work, we plan to analyze eigenvectors, which may help to identify the spatial patterns of network reorganization across the seizure states and reveal which electrode drives transitions between states, which has the potential to improve targeting for neurostimulation.
Future work will focus on developing conditions to guarantee a feasible control solution by verifying stabilizability with state-feedback. Certain network configurations that are well-characterized by fractional-order models may be fundamentally resistant to state-feedback interventions due to structural constraints, pathological states, or intrinsic nonlinearities that emerge during control attempts [62]. The numerical ill-conditioning observed in failed cases may thus reflect genuine physiological properties of these seizures rather than modeling inadequacies. The heterogeneity in control outcomes also points toward the need for patient-specific or etiology-specific controller designs. Patients with structural lesions, certain epilepsy etiologies, or particular network architectures may require alternative control strategies, such as nonlinear control, higher-order perturbations, or targeting different network nodes. Larger studies grouping patients by clinical characteristics may reveal which characteristics are more favorable to state-feedback control versus those requiring more sophisticated interventions [63, 62]. Finally, we evaluated our control framework in silico simulations using estimated models from real-world iEEG data. We plan to test our control framework in vivo and in vitro using animal models and eventually in a clinical setting, which would help us assess the real-world feasibility and efficacy of fractional-order model-based neurostimulation for seizure suppression.
@@ Methods
where ${x}[k]=(x_{1}[k],x_{2}[k],\ldots,x_{N}[k])^{{T}}$ is a state vector of the NN pre-processed iEEG recordings at discrete time step $k\in\mathbb{N}$. $A\in\mathbb{R}^{N\times N}$ is a matrix that models the spatial coupling between states. The multi-scale behavior is described using a discrete-time fractional derivative operator $\Delta^{\alpha}$, which is the Grünwald-Letnikov discretization of the fractional derivative (Chpt. 2, [17]). The vector of fractional-order exponents ${\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{N})^{\mathsf{T}}$ computes the multi-scale property for each recording, which determines how strongly each measurement depends on its past measurements. Smaller values of $\alpha$ indicate a stronger long-range temporal dependency. A value of $\alpha=1$ corresponds to a standard first-order linear time-invariant dynamical system with no multi-scale behavior. This relationship is described for each electrode $i$ by:
@@
To improve the estimation of AA and account for potential unmeasured influences, we used an iterative estimation approach [56] to estimate the parameters and unknown inputs. We augmented Eq. (1) to include an input term:
@@
where $u[k]$ represents any external inputs to the system from unmeasured sources (e.g., brain regions outside the recording area, subcortical inputs, or modeling errors). The matrix $B$ was initialized heuristically from $A^{(0)}$, which is the estimated AA matrix during the first iteration, by selecting columns that represent the most influential electrodes in the network, assuming half the recorded electrodes could receive external inputs. For each time iteration $i$, we estimated the inputs by solving a least absolute shrinkage and selection operator (LASSO) problem:
@@
The model quality was evaluated by simulating signals using the estimated parameters AA and $\alpha$. Starting from Eq. (1) and expanding the fractional derivative using the Grünwald-Letnikov weights from Eq. (3) with $\Psi(\alpha_{i},0)=1$:
@@
This controller was applied to ictal windows, starting at seizure onset. We evaluated controller effectiveness by comparing the eigenvalues of $A_{0}$ matrix (uncontrolled) and $A_{0}+\tilde{A}$ matrix (controlled) during the ictal period. Additionally, to visualize the effect of control on signal dynamics, we reconstructed signals using Eq. (9), substituting $A_{\text{ctrl}}=A+\tilde{A}$ for AA, to visualize the effect of control on signal amplitudes.